The Kuyatts' house is twice as old as the Marconis' house. Eight years ago, the Kuyatts' house was three times as old as the Marconis' house. How old is each house?
Answer by jorel1380(3719)

K=2M
K-8=3(M-8)
2M-8=3M-24
16=M
32=K
The Kuyatts' house is 32 years old. The Marconi's house is 16 years old.